OpenStudy (anonymous):

I f I have a calculator and I'm doing bar notation, how do I put the equation?

OpenStudy (anonymous):

can you be a bit more specific? are you trying, for example, to find the fraction for \[\overline{.4}\]?

OpenStudy (anonymous):

you can't put a bar in your calculator, unless i am confused about what you actually want to do

OpenStudy (anonymous):

yes but like this \[\frac{ 3 }{ 5 } \] as a decimal

OpenStudy (anonymous):

ok as a decimal \[\frac{3}{5}=0.6\] it does not repeat

OpenStudy (anonymous):

OOOOOH i see what you mean!!!!!

OpenStudy (anonymous):

you want do compute this \[\frac{3}{5}\] on a calculator right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

that is what you mean by "bar notation" yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok i was confused that is a fraction bar and \[\frac{a}{b}\] means \[a\div b\]

OpenStudy (anonymous):

so if you want \[\frac{3}{5}\] on a calculator, type in \(3\div 5\)

OpenStudy (anonymous):

is that what you were trying to do?

OpenStudy (anonymous):

Ya and they said to do with it paper and pencil you have to put \[5\div3.0\]

OpenStudy (anonymous):

So I used a calculator instead.

OpenStudy (anonymous):

what is the exact question?

OpenStudy (anonymous):

Write each fraction or mixed number as a decimal. Use bar notation if the decimal is a repeating decimal.

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

ok i got it name a fraction, and we can do it

OpenStudy (anonymous):

Ok and thank you for this

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

\[7\frac{ 9 }{ 22 }\]

OpenStudy (anonymous):

ok so this is a mixed number and you want a decimal right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

you know it is going to be \(7.something\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

to find the "something" with a calculator, type in \[9\div 22\]

OpenStudy (anonymous):

let me know what you get, then we can finish

OpenStudy (anonymous):

:O its repeating.

OpenStudy (anonymous):

yeah i know then we can use another "bar" write what you get

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[0.409 \] I cant put the notation bar

OpenStudy (anonymous):

ok but you see what repeats right? not the 4, but only the 09090909...

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so it is written like this: i am going to write it large \[\huge 7.4\overline{09}\]

OpenStudy (anonymous):

Yes how do you put the notation bar?

OpenStudy (anonymous):

that bar up top indicates that the \(09\) part is repeating

OpenStudy (anonymous):

trade secret, nsa clearance only

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

actually i am writing latex code you probably don't want to start doing that, but if you want to see the code, right click and select "show math as" then "latex"

OpenStudy (anonymous):

I have this problem \[\frac{ 17 }{ 24 }\] and I cant solve it

OpenStudy (anonymous):

what did you put in the calculator?

OpenStudy (anonymous):

i put \[17\div24\]

OpenStudy (anonymous):

ok good and what exactly did you get?

OpenStudy (anonymous):

I got 0.3541666666666666666667

OpenStudy (anonymous):

i think you must have hit the buttons wrong the question is \[\frac{17}{24}\] right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

try again then i will show you what i get

OpenStudy (anonymous):

0.70833333333333333

OpenStudy (anonymous):

yeah that looks better

OpenStudy (anonymous):

now if you want to write with a bar, put \[\large0.78\overline{3}\]

OpenStudy (anonymous):

you got this?

OpenStudy (anonymous):

il try ill message you if i need help

OpenStudy (anonymous):

ok just post, now that i know what you are doing, i can help